OpenStudy (anonymous):

What is the equation of a line passing through (2,-1) and parallel to the line represented by the equation y=2x+1?

OpenStudy (imer):

Parallel lines have same slope, use differentiation to find the slope of the equation and then create an equation for line using the point (2,-1)

OpenStudy (imer):

\[y'=2\]means slope of the parallel line is "2". \[y-y0=m(x-x0) \rightarrow y-(-1)=2(x-2) \rightarrow y=2x-5\]

OpenStudy (imer):

If you have not covered differentiation, you don't need to worry, just look for slope in the equation \[y=2x+1\] which is in form of \[y=mx+c\] where "m" is the slope, and for sure slope is "2"